Assume that the earth’s rate of rotation were increased to such an extent that the

apparent weight of an object on the equator was zero. What would then be the period

of rotation of the earth?

Radius of the earth = 6380 km


1
Expert's answer
2021-04-12T06:57:52-0400

Answer

"g^{'}=g-R\omega^2"


Given they "g^{'}=0kg\/m^2"


So

"\omega=\sqrt{\frac{g}{R}}"


"\omega=\sqrt{\frac{9.8}{6380000}}"


And time period

"T=\frac{2\pi}{\omega}=\frac{2\pi}{\sqrt{\frac{9.8}{6380000}}}=4.1\times10^6s"